The function x(t)=cos(2, where f is the sinusoidal frequency, sampled by replacing t with nT, where T is
the sampling period and n is an integer. The results is the sequence x[n]=cos(2If f=1 MHz=1,000,000, and
T is 0.1 microseconds. Write expression for x[n]. Plot x[n] over -8 < n < 8
Answer #1

Given x (t) = Cos/211 ft) where t = AMHz = 10 Hz this continous signal convert into discrete signal put t=nT a(m) = Cos (21f4 x (t) = Cos(eft) not given T then Answer below here t=nT x (M) = Cos(2fRT) - Cos(2fTn) = Cos (2x10x1x10 m) x(n) = cos(120)